You test your vertical leap and find that you can jump to a height 0.51 m above the ground.
a) What velocity do you need to have when you leave the ground to go that high?
  

b) How much hang time will you have? (How long are you off the ground?)
Answer #1

Then H=u^2/2g

T=0.319 sec

Total time =2*0.319 =0.638 sec